OpenStudy (anonymous):

y=sin(x^2)cos(x^2) find the derivative I think I'm supposed to do the product rule and I get sin(x^2)-sin(x^2)+cos(x^2)cos(x^)2 But what do I do after that?

OpenStudy (anonymous):

missing the chair rule

OpenStudy (anonymous):

*chain rule

OpenStudy (anonymous):

easiest to keep from getting bogged down in trig is to recall that \[\sin(x)\cos(x)=\frac{1}{2}\sin(2x)\]

OpenStudy (anonymous):

so you have \[\frac{1}{2}\sin(2x^2)\] take the derivative using the chain rule get \[\frac{1}{2}\cos(2x^2)\times 4x\] \[2x\cos(2x^2)\]

OpenStudy (anonymous):

otherwise using chain and product rule it is a big fat mess

OpenStudy (anonymous):

That makes it so much easier. Ya the chain rule was getting pretty busy. Thanks.